OpenStudy (anonymous):

What is the pH of a solution with a concentration of 4.7 × 10-4 molar H3O+? Show, or explain, the work used to solve this problem. Technically a math question ;D

OpenStudy (kira_yamato):

pH = -log[H3O +] pH = -log (4.7 x 10^-4) ≈ 3.33

OpenStudy (anonymous):

Thank you times a million!

OpenStudy (anonymous):

the result is 6

OpenStudy (kira_yamato):

Kumar-san, may I know how you managed to get 6? Thank you...

OpenStudy (kunal):

i too think i should be near about 2-3

OpenStudy (anonymous):

pH = -log [H+] 1. pH = - log(1.8 x 10^-4)= 3.74

OpenStudy (anonymous):

haha now I have three results! I'm sticking to @kira_yamato

OpenStudy (anonymous):

That is wrong.

OpenStudy (anonymous):

It's pH = -log [H+] 1. pH = - log(1.8 x 10^-4)= 3.74

OpenStudy (kunal):

it should be -log(4.7x10^-4) and not of 1.8x10^-4

OpenStudy (anonymous):

how?

OpenStudy (kunal):

in question it is 4.7x10^-4

OpenStudy (farmdawgnation):

This question belongs in Chem. Thanks!

OpenStudy (anonymous):

True

OpenStudy (kunal):

yaaaa.....wrong place